Solve the system by substitution
4x+3y=30
y=2x

Answer by rapaljer(4551)

4x+3y=30
y=2x

Take the y = 2x in the second equation, and substitute into the first equation:
4x + 3y = 30
4x + 3(____) = 30
4x + 3(2x)= 30
4x + 6x = 30
10x = 30
x= 3

Substitute x= 3 into the y = 2x equation (easiest!)
y= 2x
y = 2(3)
y = 6
Check in the OTHER equation (the first one!):
4x+3y = 30
4(3) + 3(6) = 30
12 + 18=30
It checks!!

R^2 at SCC